OpenStudy (anonymous):

. Find the function (do not submit a graph) that is finally graphed after the following transformations are applied to the given function. Please be sure to show all of your work. f(x)=1/x Shift down 4 units, shift left 1 unit, and reflect about the x-axis

OpenStudy (anonymous):

Just use the same method that I showed you in your previous question. \[f(x)+a\] are vertical shifts \[f(x-a)\] are horiztonal shifts \[-f(x)\] are reflections about the x-axis and \[f(-x)\] are reflections about the y-axis if you need that in a later problem. Note that you must define new functions after each operation. Sof if you wanted, for instance, to shift \[f(x)=x^2\] down 3 and then left 5, you would first shift down 3 to \[f_1(x)=x^2-3\] and then left 5 to \[f_2(x)=(x+5)^2-3\]